OpenStudy (anonymous):

what is the density of an object that is 84g and 2cm 1 cm 1cm?

OpenStudy (anonymous):

See : \[Density = \frac{Mass}{Volume}\]

OpenStudy (anonymous):

Mass = 84 g here.. And volume = 2*1*1 = 2 cm^3.

OpenStudy (anonymous):

42g 3^CM?

OpenStudy (anonymous):

Do you mean : \(42 \; \; g \; cm^3\) ??

OpenStudy (anonymous):

If YES, then you are RIGHT..

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

ty